Shares of Bleckwell Remodelers common stock are currently selling for $42.50 a share. The last annual dividend paid was $2.50 per share. The market rate of return is 15%. At what rate is the dividend growing?
Answer #1

market rate of return=(D1/Current price)+Growth rate

0.15=(2.5*(1+Growth rate)/42.5)+Growth rate

0.15=(2.5+2.5*Growth rate)/42.5+Growth rate

0.15=0.0588235294+0.0588235294*Growth rate+Growth rate

Growth rate=(0.15-0.0588235294)/(0.0588235294+1)

=8.61%(Approx).